OpenStudy (anonymous):

can any help t=3n+6n

OpenStudy (akshay_budhkar):

t=(6+3)n t=9n

OpenStudy (anonymous):

what about this t=2n-2

OpenStudy (akshay_budhkar):

it cannot be solved further or did you forget an n?

OpenStudy (anonymous):

or.. t=2(n-1)

OpenStudy (anonymous):

thats what teacher gave it to me t=2n-2

OpenStudy (akshay_budhkar):

it cannot be solved further then :)

OpenStudy (anonymous):

ic?

OpenStudy (akshay_budhkar):

ic??

OpenStudy (anonymous):

i m getting confuse .com

OpenStudy (anonymous):

thx anyway got to ask teacher again

OpenStudy (akshay_budhkar):

we cannot go further in this question t=2n-2 thats it what now?